10) What is the pH of a solution with [H3O+] = 1.0 x 10-9 M? -B)-90- C) 5.0 -D)-5.0 E) 9.0 14-9 =
Answer #1

[H3O+] = 1.0*10-9M

pH = -log[H3O+] = -log(1.0*10-9)

= -log1.0 + 9log10

= 0 + 9

pH = 9.0 (E)